How many gallons each of 25% alcohol and 35% alcohol should be mixed to obtain 20 gallons of 32% alcohol?
Answer by checkley71(8403)

.35x+.25(20-x)=.32*20
.35x+5-.25x=6.2
.1x=6.2-5
.1x=1.2
x=1.2/.1
x=12 gallons of 35 % alcohol is needed.
20-12=8 gallons of 25% alcohol is needed.
proof
.35*12+.25(20-12)=.32*20
4.2+.25*8=6.2
4.2+2=6.2
6.2=6.2